Hey! Can you please help me with the following two questions I am highly confused and I don't really understand them at all. Thank you!! 16, Please help me with them, we have a test this thursday and problems that are similar are supposed to be on there. Thank you so much!!!

1.The pH of a 0.14-M solution of phthalic acid (H2C8H4O4) is measured to be 1.91. Use this information to determine a value of Ka for phthalic acid.

  • H2C8H4O4(aq) + H2O(l) HC8H4O4-(aq) + H3O+(aq) Ka =______

2.What is the pH of a 0.142 M aqueous solution of sodium acetate, NaCH3COO? (Ka for CH3COOH = 1.8×10-5)

pH =______
Answer #1

ANSWER is Given: - Comentration of plethale and = 0.14 pH = 1.91 To find : ka=9 solution : H₂ CG H₂O4 (aq) + H₂O(e) H Ca HayTo find 0.142M aqueous solution of soduin aulat. Ka for Chaloon 1.88105 - pred solution: sodun acetate is the salt of weak anso x = N 0.142(5.55X10-10) N 7.881810-11 8.8819 x 10-6 x = 8.8819 X10-6 x = Гон1 Also, we know pH + poH = 14 pH = 14-pon = P4